# Towards an Agnostic Algorithm for Sampling Empirical Structure Models

### v1

@@ Appendix A Algorithm
We give further details regarding the implementation of the optimisation algorithm. For the readers convenience, listing 1 presenting a high-level overview of the optimisation algorithm is reproduced here.
Listing 3: The Optimisation Algorithm (reprinted)
@@
A random decreasing function is generated to serve as the starting distribution. To do this, we use a recursive subdivision method described in listing 2 which is reprinted below for the readers convenience.
Listing 4: The Binary Subdivision Algorithm (reprinted)
@@
To calculate opt_step, we determine goal functions $F_{1}$ and $F_{2}$ and a gradient. The first goal function we wish to minimise is the sum of the costs of being far away from the $d$ observed gravitational moments $J_{2n}$:
@@
The gravitational moments calculated by the Theory of Figures algorithm PyToF (Morf 2025) are denoted as $J_{2n,\text{ToF}}(\rho)$. The observed gravitational moments are denoted as $J_{2n}^{*}$ and their standard deviations as $\sigma\mkern-3.0muJ_{2n}$. For numerical stability, we include a second cost function for deviating from the planet’s mass:
@@
where $R_{m}$ and $R_{eq}$ are the mean and equatorial radius, respectively. $S_{2n,\text{ToF}}^{(-1)}$ denotes the last (innermost) and $S_{2n,\text{ToF}}^{(-2)}$ the second to last result of $S_{2n}(z)$ as defined by the Theory of Figures (see Appendix A of Morf et al. 2024) and calculated by PyToF. $\mathbf{e^{p}}$ denotes the element-wise exponential of the parameter vector. $\mathbf{{e^{p}}/{w}}$ and $\mathbf{{w}/{e^{p}}}$ are defined analogously. $\Sigma\mathbf{{e^{p}}/{w}}$ denotes the cumulative sum vector of $\mathbf{{e^{p}}/{w}}$ and $\overset{\leftarrow}{\Sigma}\mathbf{{w}/{e^{p}}}$ denotes the reverse cumulative sum vector of $\mathbf{{w}/{e^{p}}}$ (a cumulative sum starting with a sum of all elements and ending with the last element only). Lastly, T denotes the usual vector product and $\odot$ the element-wise product.
@@
Before each epoch, we check for convergence. We do this by comparing the change in cost. If the exponential moving average of the relative change in cost is close enough to 1 (difference of less than 0.0005), we terminate and return the following results:
@@ Appendix B Gradient calculation
where $S_{2n,\text{ToF}}^{(-i)}$ denotes the last/innermost (i=1i=-1) and second to last (i=2i=-2) value calculated by the Theory of Figures. The small change $\Delta z$ does not have to be further specified, it vanishes in the end result. The middle derivative (denoted element-wise here, but actually a vector) is easily calculated by its inverse, which we can approximate as
@@
since the difference between $\rho_{k}$ and $\rho_{k-1}$ is exactly this term. We calculate the leftmost derivative element-wise: Write $\partial_{i}$ for the partial derivative by $p_{i}$, therefore $\frac{\partial}{\partial p_{i}}$. The $k$-th entry of $\rho$ is $\rho_{k}=\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\alpha$. Thus:
@@
The first derivative is only non-zero if $k>i$, since $p_{i}$ only appears in the sum $\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\alpha$ for $k>i$. Therefore
$\partial_{i}\left(\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\right)=\begin{cases}\partial_{i}\frac{e^{p_{i}}}{w_{i}}=\frac{e^{p_{i}}}{w_{i}}&k>i\\ 0&k\leq i\end{cases},$ (22)
since $w_{i}$ is constant with respect to $p_{i}$. The second derivative $\partial_{i}\alpha$ is
@@
Remember $m_{calc}=4\pi\sum_{k=0}^{N-1}\tilde{\rho}_{k}l_{k}^{2}\ \Delta l$, where $\tilde{\rho}_{k}=\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}$, and note that $l_{k}$ is constant with respect to the parameters. Thus
@@ Appendix C Bias
Formally, we may call the solution space $S$. An algorithm $A$ is a random variable that produces density profiles $\rho\in S$. It has a probability density function $f_{A}:S\to\mathbb{R}_{\geq 0}$ which fulfils
@@
for any observable event $R$. The probability measure on $S$ is $\mu$. We call the algorithm $A$ unbiased if its probability density function $f_{A}$ is equal to $1$ almost everywhere. This is the formal definition of all profiles having equal likelihood of being generated. Note that this definition is stronger than just demanding $\Pr[A=\rho]$ be equal for all $\rho\in S$. Every algorithm fulfils this weaker requirement, because the probability to generate exactly $\rho$ is $\mu(\{\rho\})=0$. As a result, a more careful definition is needed: We also require that all regions $R\subseteq S$ have proportional likelihood $\int_{R}f_{A}\ \mathrm{d}\mu=\mu(R)$.
To define a concept of distance between profiles, we employ the infinity norm
@@
which allows us to approximate any region in $S$ as the countable union of distinct neighbourhoods $N_{\rho}^{\varepsilon}$ of radius $\varepsilon/2$ around $\rho$:
@@
We call two density profiles $\rho$ and $\rho^{\prime}$ similar or $\varepsilon$-close if $\rho^{\prime}\in N^{\varepsilon}_{\rho}$. From now on, we will refer to the measure of $N^{\varepsilon}_{\rho}$ as its volume, and write $V(N^{\varepsilon}_{\rho})$.
@@ C.1 Generation method
$\mkern 100.0muV(N_{\rho}^{\varepsilon})\leq\prod_{k=1}^{N-2}\left[\min\left(\frac{\varepsilon}{2},\Delta\rho_{k-1}\right)+\min\left(\frac{\varepsilon}{2},\Delta\rho_{k}\right)\right].$
Using only half of each increase ($\Delta\rho_{k-1}/2$ and $\Delta\rho_{k}/2$) guarantees no overlap between increases but thus undercounts possible neighbours, whereas using the full increase ($\Delta\rho_{k-1}$ and $\Delta\rho_{k}$) both times has overlap and thus overcounts. For the average case, it holds $\varepsilon\leq\Delta\rho_{k}$ for sufficiently small $\varepsilon$ and thus
@@
For a given $\varepsilon$ (that is, for a given level of stringency as to what two profiles are considered similar) and mall enough $\Delta\rho_{k-1},\Delta\rho_{k}<\varepsilon$, the volume becomes constant (const. $\in[0.5,1]$), which is what was desired.
@@ C.2 Gradient method
To test our results, we drew 512 successful solutions from the correlated Uranus run at random and used these as references $\rho^{*}$. To calculate the discoverability of a given $\rho^{*}$, we considered all the $n$ results $\rho^{(i)}$ that are $\varepsilon$-close to $\rho^{*}$ for a given $\varepsilon$. We then calculated the average distance between all of them
$\binom{n}{2}^{-1}\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{n}\lVert\rho^{(i)}-\rho^{(j)}\rVert_{\infty}$ (50)
@@ D.1 Discontinuity detection
Regarding the numerical detection of a Discontinuity, one could iterate over all interval sizes and find all corresponding Jumps and then find their largest intersection of Maximally Steep Jumps for each Cluster. However, this method is very slow. We thus use an approximate solution using a convolution (sliding window). An averaging sliding window (for example [0.25, 0.25, 0.25, 0.25]), would only guarantee to find Discontinuities of the size of the window. This is because smaller Discontinuities would get squashed (in our example by the factor 0.25) and larger Discontinuities could erroneously be split. For example …, 125, 125, 0, 125, 125, … should be one Discontinuity but would get seen as two with [0.5, 0.5] being the sliding window. A constant sliding window (for example [1, 1, 1, 1]) is also not usable, since it would misidentify increases consistently below the density criterion as Discontinuities, especially at large window sizes. Therefore, we used a sliding window with varying values: We chose a window size of $n=9$ and want to guarantee that we will find all Discontinuities of any size up to $n=9$. The window hence must contain one entry $\geq 1$ to find 1-Discontinuities, two entries $\geq 0.5$ to find 2-Discontinuities, and so on. $n=9$ requires nine entries $\geq 1/9$ which ultimately yields [1/8, 1/6, 1/4, 1/2, 1, 1/3, 1/5, 1/7, 1/9] as a sliding window. Since it is highly likely that Discontinuities larger than this window will have subsections within this window that are captured, this essentially guarantees all Discontinuities are found. In summary, this sliding window ensures that a Discontinuity even of only size 1 will certainly at some point be detected by exceeding the discontinuity criterion. The same holds true for a Discontinuity spread over 9 steps. However, this method is still too generous for the same reason as the above problem with the constant sliding window of [1, 1, 1, 1]. Therefore, some post-processing has to be done, where we iteratively cut potential ‘Discontinuities’ on both sides as long as the smaller ‘Discontinuity’ has better steepness and as long as the cut itself is not steeper than the discontinuity criterion. If the final resulting ‘Discontinuity’ has a total height $\geq k\cdot c_{D}$, we have successfully detected a true Discontinuity.
@@ D.2 Further results
We show the values of Figure 11 in Table 3. Figures 14 (Uranus) and 15 (Neptune) separately show the average Discontinuity height $\overline{\Delta\rho}$ and width $\overline{\Delta r}$ combined in Figure 9 at their respective average normalised radii. Note that profiles that have no Discontinuity at any given location also contribute to the averages with value zero. Artifacts from the starting model generation (see Section 4) are clearly visible in Figure 15. Finally, we show the ‘raw’ average density increase for all successful profiles in Figure 16. This Figure differs from the top panels of Figures 14 and 15 because we do not show the average Discontinuity height, but rather the general average density increase as a function of normalised radius for all successful density profiles. However, increases below the discontinuity criterion still only contribute with value zero to the average.
Figure 14: Average Discontinuity heights (top) and widths (bottom) of successful density profiles. For example, Discontinuities with an average location of $r/R=0.5$ possess an average width of $\overline{\Delta r}\sim 0.005\Delta\ell$. Profiles with no Discontinuities also contribute to the averages with value zero.

### v3

@@ Appendix A Algorithm
We give further details regarding the implementation of the optimisation algorithm. For the reader’s convenience, listing 1 presenting a high-level overview of the optimisation algorithm is reproduced here.
Listing 3: The optimisation algorithm (reprinted).
@@
A random decreasing function is generated to serve as the starting distribution. To do this, we use a recursive subdivision method described in listing 2 which is reprinted below for the reader’s convenience.
Listing 4: The binary subdivision algorithm. (reprinted)
@@
To calculate opt_step, we determine goal functions $F_{1}$ and $F_{2}$ and a gradient. The first goal function we wished to minimise was the sum of the costs of being far away from the $d$ observed gravitational moments $J_{2n}$:
@@
The gravitational moments calculated by the ToF algorithm PyToF (Morf 2025) are denoted as $J_{2n,\text{ToF}}(\rho)$. The observed gravitational moments are denoted as $J_{2n}^{*}$ and their standard deviations as $\sigma\mkern-3.0muJ_{2n}$. For numerical stability, we include a second cost function for deviating from the planet’s mass:
@@
where $R_{m}$ and $R_{eq}$ are the mean and equatorial radius, respectively. $S_{2n,\text{ToF}}^{(-1)}$ denotes the last (innermost) and $S_{2n,\text{ToF}}^{(-2)}$ the second to last result of $S_{2n}(z)$ as defined by the ToF (see Appendix A of Morf et al. 2024) and calculated by PyToF. $\mathbf{e^{p}}$ denotes the element-wise exponential of the parameter vector. $\mathbf{{e^{p}}/{w}}$ and $\mathbf{{w}/{e^{p}}}$ are defined analogously. $\Sigma\mathbf{{e^{p}}/{w}}$ denotes the cumulative sum vector of $\mathbf{{e^{p}}/{w}}$ and $\overset{\leftarrow}{\Sigma}\mathbf{{w}/{e^{p}}}$ denotes the reverse cumulative sum vector of $\mathbf{{w}/{e^{p}}}$ (a cumulative sum starting with a sum of all elements and ending with the last element only). Lastly, T denotes the usual vector product and $\odot$ the element-wise product.
@@
Before each epoch, we check for convergence. We do this by comparing the change in cost. If the exponential moving average of the relative change in cost is close enough to one (difference of less than 0.0005), we terminate and return the following results:
@@ Appendix B Gradient calculation
where $S_{2n,\text{ToF}}^{(-i)}$ denotes the last/innermost (i=1i=-1) and second to last (i=2i=-2) value calculated by the ToF. The small change $\Delta z$ does not have to be further specified, it vanishes in the end result. The middle derivative (denoted element-wise here, but actually a vector) is easily calculated by its inverse, which we can approximate as
@@
since the difference between $\rho_{k}$ and $\rho_{k-1}$ is exactly this term. We calculated the leftmost derivative element-wise: Write $\partial_{i}$ for the partial derivative by $p_{i}$, and therefore $\frac{\partial}{\partial p_{i}}$. The $k$-th entry of $\rho$ is $\rho_{k}=\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\alpha$. Thus,
@@
The first derivative is only non-zero if $k>i$, since $p_{i}$ only appears in the sum $\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\alpha$ for $k>i$. Therefore,
$\partial_{i}\left(\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}\right)=\begin{cases}\partial_{i}\frac{e^{p_{i}}}{w_{i}}=\frac{e^{p_{i}}}{w_{i}}&k>i\\ 0&k\leq i\end{cases}$ (22)
since $w_{i}$ is constant with respect to $p_{i}$. The second derivative, $\partial_{i}\alpha$, is
@@
We remind that $m_{calc}=4\pi\sum_{k=0}^{N-1}\tilde{\rho}_{k}l_{k}^{2}\ \Delta l$, where $\tilde{\rho}_{k}=\sum_{j=0}^{k-1}\frac{e^{p_{j}}}{w_{j}}$ and note that $l_{k}$ is constant with respect to the parameters. Thus,
@@ Appendix C Bias
Formally, we may call the solution space $S$. An algorithm $A$ is a random variable that produces density profiles $\rho\in S$. It has a probability density function $f_{A}:S\to\mathbb{R}_{\geq 0}$, which fulfils
@@
for any observable event, $R$. The probability measure on $S$ is $\mu$. We call the algorithm $A$ unbiased if its probability density function $f_{A}$ is equal to $1$ almost everywhere. This is the formal definition of all profiles having equal likelihood of being generated. Note that this definition is stronger than just demanding $\Pr[A=\rho]$ be equal for all $\rho\in S$. Every algorithm fulfils this weaker requirement, because the probability to generate exactly $\rho$ is $\mu(\{\rho\})=0$. As a result, a more careful definition is needed: We also require that all regions $R\subseteq S$ have proportional likelihood $\int_{R}f_{A}\ \mathrm{d}\mu=\mu(R)$.
To define a concept of distance between profiles, we employed the infinity norm
@@
which allowed us to approximate any region in $S$ as the countable union of distinct neighbourhoods $N_{\rho}^{\varepsilon}$ of radius $\varepsilon/2$ around $\rho$:
@@
We call two density profiles, $\rho$ and $\rho^{\prime}$, similar or $\varepsilon$-close if $\rho^{\prime}\in N^{\varepsilon}_{\rho}$. From now on, we refer to the measure of $N^{\varepsilon}_{\rho}$ as its volume, and write $V(N^{\varepsilon}_{\rho})$.
@@ C.1 Generation method
$\mkern 100.0muV(N_{\rho}^{\varepsilon})\leq\prod_{k=1}^{N-2}\left[\min\left(\frac{\varepsilon}{2},\Delta\rho_{k-1}\right)+\min\left(\frac{\varepsilon}{2},\Delta\rho_{k}\right)\right],$
which is a bound tighter (and hence better) compared to a bound relying on $\varepsilon$ alone. Using only half of each increase ($\Delta\rho_{k-1}/2$ and $\Delta\rho_{k}/2$) guarantees no overlap between increases but thus undercounts possible neighbours, whereas using the full increase ($\Delta\rho_{k-1}$ and $\Delta\rho_{k}$) both times has overlap and thus overcounts. For the average case, it holds $\varepsilon\leq\Delta\rho_{k}$ for sufficiently small $\varepsilon$, and thus
@@
For a given $\varepsilon$ (that is, for a given level of stringency as to what two profiles are considered similar) and small enough $\Delta\rho_{k-1},\Delta\rho_{k}<\varepsilon$, the volume becomes constant (const. $\in[0.5,1]$), which is what was desired.
@@ C.2 Gradient method
To test our results, we drew 512 successful solutions from the correlated Uranus run at random and used these as references $\rho^{*}$. To calculate the discoverability of a given $\rho^{*}$, we considered all the $n$ results $\rho^{(i)}$ that are $\varepsilon$-close to $\rho^{*}$ for a given $\varepsilon$. We then calculated the average distance between all of them:
$\binom{n}{2}^{-1}\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{n}\lVert\rho^{(i)}-\rho^{(j)}\rVert_{\infty},$ (50)
@@ D.1 Discontinuity detection
Regarding the numerical detection of a discontinuity, one could iterate over all interval sizes and find all corresponding jumps and then find their largest intersection of maximally steep jumps for each cluster. However, this method is very slow. We thus use an approximate solution using a convolution (sliding window). An averaging sliding window (for example [0.25, 0.25, 0.25, 0.25]), would only guarantee to find discontinuities of the size of the window. This is because smaller discontinuities would get squashed (in our example by the factor 0.25) and larger discontinuities could erroneously be split. For example …, 125, 125, 0, 125, 125, … should be one discontinuity but would get identified as two with [0.5, 0.5] being the sliding window. A constant sliding window (for example [1, 1, 1, 1]) is also not usable, since it would misidentify increases consistently below the density criterion as discontinuities, especially at large window sizes. Therefore, we used a sliding window with varying values: We chose a window size of $n=9$ and want to guarantee that we will find all discontinuities of any size up to $n=9$. The window hence must contain one entry $\geq 1$ to find discontinuities of length one, two entries $\geq 0.5$ to find discontinuities of length two, and so on. $n=9$ requires nine entries $\geq 1/9$ which ultimately yields [1/8, 1/6, 1/4, 1/2, 1, 1/3, 1/5, 1/7, 1/9] as a sliding window. Since it is highly likely that discontinuities larger than this window will have subsections within this window that are captured, this essentially guarantees all discontinuities are found. In summary, this sliding window ensures that a discontinuity even of only size 1 will certainly at some point be detected by exceeding the discontinuity criterion. The same holds true for a discontinuity spread over nine steps. However, this method is still too generous for the same reason as the above problem with the constant sliding window of [1, 1, 1, 1]. Therefore, some post-processing has to be done, where we iteratively cut potential ‘discontinuities’ on both sides as long as the smaller ‘discontinuity’ has better steepness and as long as the cut itself is not steeper than the discontinuity criterion. If the final resulting ‘discontinuity’ has a total height $\geq k\cdot c_{D}$, we have successfully detected a true discontinuity.
@@ D.2 Further results
We show the values of Figure 11 in Table 3. Figures 14 (Uranus) and 15 (Neptune) separately show the average discontinuity height $\overline{\Delta\rho}$ and width $\overline{\Delta r}$ combined in Figure 9 at their respective average normalised radii. Note that profiles that have no discontinuity at any given location also contribute to the averages with value zero. Artefacts from the starting model generation (see Section 4) are clearly visible in Figure 15. Finally, we show the ‘raw’ average density increase for all successful profiles in Figure 16. This Figure differs from the left panels of Figures 14 and 15 because we do not show the average discontinuity height, but rather the general average density increase as a function of normalised radius for all successful density profiles. However, increases below the discontinuity criterion still only contribute with value zero to the average.
Figure 14: Average discontinuity heights (left) and widths (right) of successful density profiles. For example, discontinuities with an average location of $r/R=0.5$ possess an average width of $\overline{\Delta r}\sim 0.005\Delta\ell$. Profiles with no discontinuities also contribute to the averages with value zero.